OpenStudy (anonymous):

how do you do (f/g) f=(x/x+1) g=x^2 -1

OpenStudy (anonymous):

Given: f(x) and g(x) then (f/g)(x)=f(x)/g(x) So you have: \[\frac{f(x)}{g(x)}=(\frac{f}{g})(x)=\frac{\frac{x}{x+1}}{x^2-1}=\frac{x}{(x+1)(x^2-1)}=\frac{x}{(x+1)(x+1)(x-1)}=\frac{x}{(x+1)^2(x-1)}\]